# Covering Multiple Objectives with a Small Set of Solutions Using Bayesian Optimization

### v1

@@ 4.4 Ablation Study
Figure 6: Ablation study comparing MOCOBO to optimization performance where a known “good” partitioning of the $T$ objectives into $K$ subsets is available in advance. We individually optimize $K$ solutions, one for each partition.
@@
With such an “oracle” partitioning in hand ahead of time, we can efficiently optimize by running $K$ independent optimization runs: one to find a single solution for each subset in the fixed, given partition. With a fixed partition, the coverage score in (1) reduces to optimizing the sum of objectives in each relevant subset independently.
In Figure 6 we compare this strategy directly to MOCOBO on three optimization tasks. We plot the best coverage score obtained by $K$ independent runs of TuRBO/LOL-BO on the fixed oracle $K$-partitioning of the objectives by combining their solutions into a single set. MOCOBO achieves the same average best coverage scores with minimal loss in optimization efficiency despite having to discover an efficient partitioning during optimization.
@@ 5 Discussion
By bridging the gap between traditional multi-objective optimization and practical coverage requirements, MOCOBO offers an effective approach to tackle critical problems in drug design and beyond. This framework extends the reach of Bayesian optimization into new domains, providing a robust solution to the challenges posed by extreme trade-offs and the need for specialized, collaborative solutions.
Impact Statement
This research includes applications in molecule and peptide design. While AI-driven biological design holds great promise for benefiting society, it is crucial to acknowledge its dual-use potential. Specifically, AI techniques designed for drug discovery could be misused to create harmful biological agents (Urbina et al., 2020).
@@ A.1 Methods for Obtaining In Vitro Minimal Inhibitory Concentration (MIC) Data
The in vitro antimicrobial activity of the peptides was assessed by using the broth microdilution assay (Cesaro et al., 2022). Minimal inhibitory concentration (MIC) values of the peptides were determined with an initial inoculum of $2\times 10^{6}$ cells $mL^{-1}$ in LB in microtiter $96$-well flat-bottom transparent plates. Aqueous solutions of the peptides were added to the plate at concentrations ranging from $0.0625$ to $64$ $\mu molL^{-1}$. The lowest concentration of peptide that inhibited 100 percent of the visible growth of bacteria was established as the MIC value in an experiment of $20h$ of exposure at 37 degrees Celsius. The optical density of the plates was measured at $600nm$ using a spectrophotometer. All assays were done as three biological replicates.
@@ A.2 Additional Results
In Figure A.1, we provide In vitro results for the two best “template free” and two best “template constrained” runs of MOCOBO for the peptide design task. Here, “best” means runs that achieved highest coverage scores according to the APEX 1.1 model. Figure A.1 provides in vitro MICs for each of the $K=4$ peptides found by each of these runs of MOCOBO, for each of the $11$ target pathogenic bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
@@
Table A.1: The best set of $K=4$ peptide sequences found by one run of MOCOBO for the “template free” peptide design task described in Section 4.2. For each of the four sequences, we provide the MIC according to the APEX model for each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.5. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TC1” in Figure 4 and Figure A.1 for in vitro MICs for this set of $K=4$ peptides.
In Table A.1, we provide an example of a covering set of peptides found by MOCOBO for the “template constrained” variation of the peptide design task.
In Figure A.2, we provide the original hdr image (leftmost panel), and the $K=4$ images produced using the $K=4$ solutions found by a single run of MOCOBO for the church image variation of the image tone mapping task.
@@
In Figure A.3 and Figure A.4, we provide examples of a covering set of trajectories found for the $T=8$ and $T=12$ variations of the rover task respectively. In each figure, a panel is shown for each of the $T$ obstacles courses, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal” end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. Each panel also shows the best among the $K$ trajectories found by a run of MOCOBO for navigating each obstacle course. An analogous plot for the $T=4$ variation of the rover optimization task is provided in the main text in Figure 5.
Figure A.2: (Leftmost Panel) Initial hdr church image for the church test image variation of the “image tone mapping” task. (Four Rightmost Panels) Images obtained by transforming the hdr church image using the best covering set of $K=4$ solutions found by a single run of MOCOBO.
@@
Table A.2: The best set of $K=3$ molecules found by one run of MOCOBO for the Ranolazine MPO multiple element molecule design task described in Section 4.2. For each of the three molecules, we provide the objective value obtained for each of the $T=6$ objectives. The best/highest objective value is in bold in each column. Each objective aims to add a different target element to Ranolazine (F, Cl, Br, Se, S, and P). The $T=6$ target elements are in bold in the smiles string (Weininger, 1988) representation of each molecule.
@@
Figure A.3: The 8 panels depict the 8 obstacle courses that the rover must navigate for the $T=8$ variation of the rover task, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal” end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating the obstacle course from among the $K=2$ covering trajectories found by a single run of MOCOBO. The first trajectory in the covering set is shown in magenta and successfully navigates obstacle courses 5, 6, 7, and 8 (Bottom Row). The second is shown in blue and successfully navigates obstacle courses 1, 2, 3, and 4 (Top Row).
@@
Figure A.4: The 12 panels depict the 12 obstacle courses that the rover must navigate for the $T=12$ variation of the rover task, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal” end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating each obstacle course from among the $K=4$ covering trajectories found by a run of MOCOBO. The first trajectory in the covering set is shown in magenta and successfully navigates obstacle courses 6 and 7 (Middle Row: Center Left and Center Right Panels). The second is shown in blue and successfully navigates obstacle courses 2 and 8 (Top Row: Center Left Panel, Middle Row: Rightmost Panel). The third is shown in purple and successfully navigates obstacle courses 4, 5, and 11 (Top Row: Rightmost Panel, Middle Row: Leftmost Panel, and Bottom Row: Center Right Panel). The fourth trajectory is shown in orange and successfully navigates obstacle courses 1, 3, 9, 10, and 12 (Top Row: Leftmost Panel and Center Right Panel, Bottom Row: Leftmost, Center Left, and Rightmost Panels).
@@ A.3 Approximation Strategy for Batch Acquisition with Expected Coverage Improvement (ECI)
In batch acquisition, we select a batch of $q>1$ candidates from each trust region for evaluation. In Equation 7, we define q-ECI, a natural extension of the ECI acquisition function defined in Equation 5 to the batch acquisition setting. q-ECI gives the expected improvement in the coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. When using batch acquisition, we aim to select a batch of $q$ points that maximize q-ECI.
We will first discuss how one would estimate q-ECI using a Monte Carlo (MC) approximation. To select a batch candidates $\hat{\mathbf{X}}_{k}$ from $\mathcal{T}_{k}$, we sample $m$ batches of $q$ points $B_{k}=\left\{\mathbf{B}_{k1},\mathbf{B}_{k2},...,\mathbf{B}_{km}\right\}$ from $\mathcal{T}_{k}$. Here $\mathbf{B}_{kj}$ is a batch of $q$ sampled points $\mathbf{B}_{kj}=\left\{\mathbf{b}_{kj1},\ldots,\mathbf{b}_{kjq}\right\}$. For each batch $\mathbf{B}_{kj}$, we sample a realization $\hat{\mathbf{Y}}_{kj}=\left\{\hat{\mathbf{y}}_{kj1},\ldots,\hat{\mathbf{y}}_{kjq}\right\}$ from the GP surrogate model posterior. We leverage these samples to compute an MC approximation to the q-ECI of each $\mathbf{B}_{kj}$:
$q\text{-CI}(\mathbf{B}_{kj})$ $=\max(0,\max_{r=1,\dots,q}c(S^{*}_{D_{s}\cup\left\{(\mathbf{b}_{kjr},\hat{\mathbf{y}}_{kjr})\right\}})-c(S^{*}_{D_{s}}))$ (8)
@@
Instead of sampling $m$ batches of candidates, we sample $m$ individual data points $P_{k}=\left\{\mathbf{p}_{k1},\mathbf{p}_{k2},...,\mathbf{p}_{km}\right\}$ from trust region $\mathcal{T}_{k}$. For each sampled point $\mathbf{p}_{kj}$, we sample a realization $\hat{\mathbf{y}}_{kj}=(\hat{f}_{1}(\mathbf{p}_{kj}),\ldots,\hat{f}_{T}(\mathbf{p}_{kj}))$. As in Section 3.1.1, we use the sampled realizations $\hat{\mathbf{y}}_{kj}$ to compute an approximate coverage improvement $CI(\mathbf{p}_{kj})$ as defined in Equation 6 for each point $\mathbf{p}_{kj}$. To obtain a batch of $q$ candidates, we then greedily select the the $q$ points $\mathbf{p}_{kj}\in P_{k}$ with the $q$ largest expected coverage improvements.
A.4 Additional Implementation Details

### v2

@@ 4.3 Ablation study
With such an “oracle” partitioning in hand ahead of time, we can efficiently optimize by running $K$ independent optimization runs: one to find a single solution for each subset in the fixed, given partition. With a fixed partition, the coverage score in (1) reduces to optimizing the sum of objectives in each relevant subset independently. In Figure 3 right, we compare this strategy directly to MOCOBO on three optimization tasks. We plot the best coverage score obtained by $K$ independent runs of TuRBO/LOL-BO on the fixed oracle $K$-partitioning of the objectives by combining their solutions into a single set. MOCOBO achieves the same average best coverage scores with minimal loss in optimization efficiency despite having to discover an efficient partitioning during optimization. See Section A.2.2 for an additional ablation study of MOCOBO when all objectives are highly conflicting.
@@ 5 Discussion
By bridging the gap between traditional multi-objective Bayesian optimization (BO) and practical coverage requirements, MOCOBO offers an effective approach to tackle critical problems in drug design and beyond. This framework extends the reach of Bayesian optimization into new domains, providing a robust solution to the challenges posed by extreme trade-offs and the need for specialized, collaborative solutions. See Section A.4 for additional discussion and limitations of MOCOBO.
Acknowledgments and Disclosure of Funding
N. Maus was supported by the National Science Foundation Graduate Research Fellowship; K. Kim was supported by a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI; J. R. Gardner was supported by NSF awards IIS-2145644 and DBI-2400135; C. de la Fuente-Nunez was supported by NIH grant R35GM138201, and by Defense Threat Reduction Agency grants HDTRA11810041, HDTRA1-21-1-0014, and HDTRA1-23-1-0001.
@@ A.1 Methods for obtaining in vitro minimal inhibitory concentration (MIC) data
The in vitro antimicrobial activity of the peptides was assessed by using the broth microdilution assay [8]. Minimal inhibitory concentration (MIC) values of the peptides were determined with an initial inoculum of $2\times 10^{6}$ cells $mL^{-1}$ in LB in microtiter $96$-well flat-bottom transparent plates. Aqueous solutions of the peptides were added to the plate at concentrations ranging from $0.0625$ to $64$ $\mu molL^{-1}$. The lowest concentration of peptide that inhibited 100 percent of the visible growth of bacteria was established as the MIC value in an experiment of $20h$ of exposure at 37 degrees Celsius. The optical density of the plates was measured at $600nm$ using a spectrophotometer. All assays were done as three biological replicates.
@@ A.2 Additional results
In this section, we provide all additional empirical results not included in the main text.
A.2.1 Additional covering sets of solutions found by MOCOBO
@@
In Figure A.1, we provide In vitro results for the two best “template free" and two best “template constrained" runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. Figure A.1 provides in vitro MICs for each of the $K=4$ peptides found by each of these runs of MOCOBO, for each of the $11$ target pathogenic bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
@@
Table A.1: The best set of $K=4$ peptide sequences found by one run of MOCOBO for the “template free" peptide design task described in Section 4.1. For each of the four sequences, we provide the MIC according to the APEX model for each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TC1" in Figure 3 left, and Figure A.1 for in vitro MICs for this set of $K=4$ peptides.
In Table A.1, we provide an example of a covering set of peptides found by MOCOBO for the “template constrained" variation of the peptide design task.
In Figure A.2, we provide the original hdr image (leftmost panel), and the $K=4$ images produced using the $K=4$ solutions found by a single run of MOCOBO for the church image variation of the image tone mapping task. An analogous result for the desk image variation of the image tone mapping task can be found in the main text in Figure 4 right. A notable apparent limitation of the existing IAA/IQA metrics we used for the image tone mapping tasks is that they all favored monochromatic images for both the church and desk images. As such, the tone-mapped images obtained appear less colorful than their hand-tuned counterparts reported in prior work.
@@
In Figure A.3 and Figure A.4, we provide examples of a covering set of trajectories found for the $T=8$ and $T=12$ variations of the rover task respectively. In each figure, a panel is shown for each of the $T$ obstacles courses, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. Each panel also shows the best among the $K$ trajectories found by a run of MOCOBO for navigating each obstacle course. An analogous plot for the $T=4$ variation of the rover optimization task is provided in the main text in Figure 4.
Figure A.2: (Leftmost Panel) Image obtained by naively compressing the dynamic range of the HDR church image in the “image tone mapping" task. (Four Rightmost Panels) Images obtained by applying tone mapping to the church image using the best covering set of $K=4$ solutions found by a single run of MOCOBO.
@@
Table A.2: The best set of $K=3$ molecules found by one run of MOCOBO for the Ranolazine MPO multiple element molecule design task described in Section 4.1. For each of the three molecules, we provide the objective value obtained for each of the $T=6$ objectives. The best/highest objective value is in bold in each column. Each objective aims to add a different target element to Ranolazine (F, Cl, Br, Se, S, and P). The $T=6$ target elements are in bold in the smiles string [69] representation of each molecule.
@@
Figure A.3: The 8 panels depict the 8 obstacle courses that the rover must navigate for the $T=8$ variation of the rover task, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating the obstacle course from among the $K=2$ covering trajectories found by a single run of MOCOBO. The first trajectory in the covering set is shown in magenta and successfully navigates obstacle courses 5, 6, 7, and 8 (Bottom Row). The second is shown in blue and successfully navigates obstacle courses 1, 2, 3, and 4 (Top Row).
@@
Figure A.4: The 12 panels depict the 12 obstacle courses that the rover must navigate for the $T=12$ variation of the rover task, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating each obstacle course from among the $K=4$ covering trajectories found by a run of MOCOBO. The first trajectory in the covering set is shown in magenta and successfully navigates obstacle courses 6 and 7 (Middle Row: Center Left and Center Right Panels). The second is shown in blue and successfully navigates obstacle courses 2 and 8 (Top Row: Center Left Panel, Middle Row: Rightmost Panel). The third is shown in purple and successfully navigates obstacle courses 4, 5, and 11 (Top Row: Rightmost Panel, Middle Row: Leftmost Panel, and Bottom Row: Center Right Panel). The fourth trajectory is shown in orange and successfully navigates obstacle courses 1, 3, 9, 10, and 12 (Top Row: Leftmost Panel and Center Right Panel, Bottom Row: Leftmost, Center Left, and Rightmost Panels).
A.2.2 Additional ablation study: $T$ highly conflicting objectives
MOCOBO is designed for cases where some, but not all, of the objectives are highly conflicting, and this is the case for all of the tasks considered in Section 4. The conflicting pairs of the objectives prevent a single solution from optimizing all objectives well. The fact that not all pairs of objectives are completely conflicting, is the reason why it is possible to cover all T objectives with a small set of K solutions. If all pairs of the T objectives are highly conflicting, there is by definition no possible set of K<T solutions such that all objectives are well optimized. In this case, it’s best to use T individually optimized solutions if your goal is to find at least one solution that well optimizes each objective. However, it is still interesting to consider the performance of MOCOBO in the setting of T objectives that are pairwise highly conflicting. To investigate this, we construct a new variation of the multi-objective rover task described in Section 4.1. We design T=3 obstacle courses such that it is impossible for the rover to take any single path that avoids all obstacles in any pair of the obstacle courses. This results in T=3 objectives that are pairwise highly conflicting. We then run 20 replications of MOCOBO on this problem with K=2, asking MOCOBO to design K=2 solutions that cover the T=3 pairwise highly conflicting objectives. For this task, MOCOBO got an average coverage score of $2.932$, while T individually optimized solutions got an average coverage score of $12.899$. It is unsurprising that T individually optimized solutions achieved better coverage here since all three objectives are pairwise highly conflicting. However, this result demonstrates that MOCOBO is able to design sets of K=2 solutions that still achieve fairly high coverage scores.
Figure A.5: The three panels depict the T=3 obstacle courses that the rover must navigate for the ablation variation of the rover task where all T=3 objectives pairwise highly conflicting. Obstacles are colored in red. The three obstacle courses are designed such that any single path that successfully avoids all obstacles in one obstacle course, must hit some obstacles in both of the other obstacle courses. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating the obstacle course from among the K=2 covering trajectories found by a single run of MOCOBO. The first trajectory in the covering set is shown in magenta and is the best trajectory for navigating obstacle courses 2 and 3 (Middle Panel and Rightmost Panels). The second is shown in blue and successfully navigates obstacle course 1 (Leftmost Panel).
Figure A.5 provides a diagram of the T=3 pairwise highly conflicting obstacle courses, and an example of one of the covering sets of K=2 trajectories found by a single run of MOCOBO. One optimized trajectory (shown in blue) is designed by MOCOBO such that it specializes to navigate the first obstacle course, navigating it without hitting any obstacles (Leftmost Panel). The other optimized trajectory (shown in magenta) is designed by MOCOBO such that it avoids all obstacles in the third obstacle course (Rightmost Panel), while also minimizing total amount of impact with the obstacles in the second obstacle course (Middle Panel). Since it is by-design impossible to avoid all obstacles in all T=3 obstacle courses with only K=2 solutions, Figure A.5 demonstrates that MOCOBO was able to successfully balance trade-offs among the objectives, designing a set of K=2 solutions that minimized the total amount of obstacle impact across the three obstacle courses.
A.2.3 Wall-clock runtime of Algorithm 1
In Table A.3, we provide an empirical evaluation of the average wall-clock runtime of Algorithm 1 with different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). We provide average run times with $N=20,200,2000,20000,200000,$ and $2000000$ for each combination of $K$ and $T$ that we ran MOCOBO on in Section 4. To improve the speed of Algorithm 1, in all experiments, we compute the inner for loop (lines 5-6 in Algorithm 1) in parallel. This involves simply computing the marginal coverage of adding all $N$ data points at once in parallel, and then updating $A$ with the point that achieved maximum marginal converge improvement. Parallel computation of the marginal coverage improvement of the $N$ points is straightforward to implement using PyTorch. In Table A.3, the “Runtime" column gives the average runtime of Algorithm 1 with this parallel computation of the of the inner for loop, and the “Runtime No Parallelization" column gives the average runtime without the parallel computation. Notice that as $N$ grows large, this parallel computation becomes essential to achieve the reasonably fast run times of Algorithm 1 needed to run MOCOBO. Average run times reported in Table A.3 are computed by running Algorithm 1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs.
N K T Runtime No Parallelization (seconds) Runtime (seconds)
20 2 4 $0.00605\pm 0.00385$ $0.00864\pm 0.00795$
200 2 4 $0.0230\pm 0.00391$ $0.00901\pm 0.00828$
2000 2 4 $0.194\pm 0.00389$ $0.00850\pm 0.00781$
20000 2 4 $2.0523\pm 0.0259$ $0.00852\pm 0.00780$
200000 2 4 $20.845\pm 0.0446$ $0.00872\pm 0.00796$
2000000 2 4 $195.740\pm 0.114$ $0.00923\pm 0.00792$
20 2 8 $0.00596\pm 0.00375$ $0.00906\pm 0.00833$
200 2 8 $0.0269\pm 0.00446$ $0.00894\pm 0.00808$
2000 2 8 $0.197\pm 0.00386$ $0.00865\pm 0.00794$
20000 2 8 $1.947\pm 0.00348$ $0.00864\pm 0.00789$
200000 2 8 $19.508\pm 0.00964$ $0.00915\pm 0.00825$
2000000 2 8 $193.317\pm 0.218$ $0.00962\pm 0.00785$
20 4 12 $0.00861\pm 0.00430$ $0.00873\pm 0.00785$
200 4 12 $0.0440\pm 0.00371$ $0.00875\pm 0.00789$
2000 4 12 $0.386\pm 0.00642$ $0.00910\pm 0.00821$
20000 4 12 $4.303\pm 0.120$ $0.00869\pm 0.00776$
200000 4 12 $38.596\pm 0.0895$ $0.00889\pm 0.00784$
2000000 4 12 $395.325\pm 0.169$ $0.0111\pm 0.00772$
20 3 6 $0.00704\pm 0.00384$ $0.00887\pm 0.00809$
200 3 6 $0.0334\pm 0.00372$ $0.00840\pm 0.00764$
2000 3 6 $0.314\pm 0.00654$ $0.00897\pm 0.00817$
20000 3 6 $2.942\pm 0.00478$ $0.00872\pm 0.00785$
200000 3 6 $28.732\pm 0.0156$ $0.00862\pm 0.00775$
2000000 3 6 $312.825\pm 0.203$ $0.00985\pm 0.00799$
20 4 7 $0.00765\pm 0.00370$ $0.00880\pm 0.00793$
200 4 7 $0.0448\pm 0.00365$ $0.00847\pm 0.00760$
2000 4 7 $0.412\pm 0.00252$ $0.00887\pm 0.00798$
20000 4 7 $3.917\pm 0.00461$ $0.00882\pm 0.00789$
200000 4 7 $38.932\pm 0.0402$ $0.00889\pm 0.00791$
2000000 4 7 $389.272\pm 0.588$ $0.0105\pm 0.00809$
20 4 11 $0.00772\pm 0.00377$ $0.00955\pm 0.00853$
200 4 11 $0.0422\pm 0.00383$ $0.00859\pm 0.00775$
2000 4 11 $0.387\pm 0.00382$ $0.00931\pm 0.00837$
20000 4 11 $4.155\pm 0.00758$ $0.00887\pm 0.00792$
200000 4 11 $41.869\pm 0.0203$ $0.00880\pm 0.00777$
2000000 4 11 $381.313\pm 0.199$ $0.0114\pm 0.00803$
Table A.3: Average wall-clock runtime of Algorithm 1 for different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). The “Runtime" column gives the average runtime of Algorithm 1 with parallel computation of the of the marginal coverage of adding the $N$ data points (parallel computation of lines 5-6). The “Runtime No Parallelization" column gives the average runtime of Algorithm 1 without this parallelization of the inner for loop. Average run times are computed by running Algorithm 1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs. Standard errors over the $10$ runs are also provided.
Efficiency of Algorithm 1
Algorithm 1 could also be made more efficient by pruning the points in $D_{s}$ with zero marginal coverage improvement after each outer-loop, since these points will continue to have zero marginal coverage improvement on subsequent loops. In future work, we plan to explore this and any other tricks that might allow us to further improve the efficiency of Algorithm 1.
@@ A.3 Approximation strategy for batch acquisition with Expected Coverage Improvement (ECI)
In batch acquisition, we select a batch of $q>1$ candidates for evaluation. In Equation 6, we define q-ECI, a natural extension of the ECI acquisition function defined in Equation 4 to the batch acquisition setting. q-ECI gives the expected improvement in the coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. When using batch acquisition, we aim to select a batch of $q$ points that maximize q-ECI.
We will first discuss how one would estimate q-ECI using a Monte Carlo (MC) approximation. To select a batch candidates $\hat{\mathbf{X}}$, we sample $m$ batches of $q$ points $B=\left\{\mathbf{B}_{1},\mathbf{B}_{2},...,\mathbf{B}_{m}\right\}$. Here $\mathbf{B}_{j}$ is a batch of $q$ sampled points $\mathbf{B}_{j}=\left\{\mathbf{b}_{j1},\ldots,\mathbf{b}_{jq}\right\}$. For each batch $\mathbf{B}_{j}$, we sample a realization $\hat{\mathbf{Y}}_{j}=\left\{\hat{\mathbf{y}}_{j1},\ldots,\hat{\mathbf{y}}_{jq}\right\}$ from the GP surrogate model posterior. We leverage these samples to compute an MC approximation to the q-ECI of each $\mathbf{B}_{j}$:
$q\text{-CI}(\mathbf{B}_{j})$ $=\max(0,\max_{r=1,\dots,q}c(S^{*}_{D_{s}\cup\left\{(\mathbf{b}_{jr},\hat{\mathbf{y}}_{jr})\right\}})-c(S^{*}_{D_{s}}))$ (7)
@@
Instead of sampling $m$ batches of candidates, we sample $m$ individual data points $P=\left\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{m}\right\}$. For each sampled point $\mathbf{p}_{j}$, we sample a realization $\hat{\mathbf{y}}_{j}=(\hat{f}_{1}(\mathbf{p}_{j}),\ldots,\hat{f}_{T}(\mathbf{p}_{j}))$. As in Section 3.1.1, we use the sampled realizations $\hat{\mathbf{y}}_{j}$ to compute an approximate coverage improvement $CI(\mathbf{p}_{j})$ as defined in Equation 5 for each point $\mathbf{p}_{j}$. To obtain a batch of $q$ candidates, we then greedily select the $q$ points $\mathbf{p}_{j}\in P$ with the $q$ largest expected coverage improvements. Note that this strategy does not involve sequential optimization of the $q$ points, as the batch of $q$ points is selected simultaneously as the points with the top-$q$ expected coverage improvements.
A.4 Limitations and future work
A primary limitation of MOCOBO is that the choice of $K$ (the covering set size) may not always be straightforward. For example, as we mention in Section 4, it may require domain knowledge to choose $K<T$ that is large enough that achieving good coverage is possible despite multiple conflicting objectives. In many practical applications, we prefer the smallest possible set size $K$ such that good coverage can still be achieved. In future work, we plan to explore methods for simultaneously optimizing both the solutions in the covering set, and the size of the covering set, balancing the trade-off between minimizing the number of solutions needed and maximizing overall coverage.
A.5 Broader impacts
This research includes applications in molecule and peptide design. While AI-driven biological design holds great promise for benefiting society, it is crucial to acknowledge its dual-use potential. Specifically, AI techniques designed for drug discovery could be misused to create harmful biological agents [65].
A.6 Compute resources
In this section, we provide all details about compute resources used to produce all results in this paper.
Table A.4: Setup of internal cluster used to run experiments.
Type Specifications
System Topology 20 nodes with 2 sockets each with 24 logical threads (total 48 threads)
Processor 1 Intel Xeon Silver 4310, 2.1 GHz (maximum 3.3 GHz) per socket
Cache 1.1 MiB L1, 30 MiB L2, and 36 MiB L3
Memory 250 GiB RAM
Accelerator 1 NVIDIA RTX A5000 per node, 2 GHZ, 24GB RAM
Compute specifications (type and memory).
We use GPU works to run all experiments and produce all empirical results provided in this paper. A single GPU was used per run of each method compared on each task. Each each run uses approximately 12-18 GB of the GPU memory. Most experiments were executed on our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table A.4). We also used cloud compute resources for two weeks to complete additional replications of some experiments. We used a total of eight RTX 4090 GPU workers from runpod.io, each with approximately 24 GB of GPU memory.
Execution time.
For the relatively inexpensive rover task, each optimization run takes approximately 1 day of execution time. For all other tasks considered, each optimization run takes approximately 3 days of execution time. To create all coverage optimization plots, we ran all methods compared $20$ times each. Completing all of the runs needed to produce all of the results in this paper required roughly $64000$ total GPU hours.
Compute resources for preliminary experiments.
Preliminary experiments refer to the initial experiments for e.g. method development that are not included as results in the paper. All preliminary experiments were done our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table A.4). We spent approximately $2000$ hours of GPU time on preliminary experiments.
A.7 Additional implementation details